# Chemical Abundances in the Leiptr Stellar Stream: a Disrupted Ultra-faint Dwarf Galaxy?

### v1

@@ 4.6 Neutron-capture elements (Sr, Ba, Eu)
Figure 11 shows that Leiptr has very low neutron-capture abundances, [Sr/Fe] [Ba/Fe] . The four Leiptr member stars have significantly lower [Sr/Fe] and [Ba/Fe] than halo stars, most dwarf galaxy stars, or any of the Milky Way globular clusters. For example, globular clusters M92 and NGC 2419 have higher [Sr/Fe] between and 0, and Leiptr’s [Sr/Fe] only falls within the extreme tail end of the halo scatter in Figure 11. The systems with the most similar neutron-capture element abundances, especially for strontium, are the ultra-faint dwarf galaxies with , which typically have low [Sr/Fe] and [Ba/Fe] between and (Frebel & Norris, 2015; Kirby et al., 2017; Ji et al., 2019).
We detect [Eu/Fe] in the two cooler Leiptr members at quite low values, . These two stars have [Ba/Eu] compared to a pure -process ratio of (e.g., Simmerer et al., 2004), suggesting the neutron-capture element abundances in Leiptr are dominated by production via the -process but potentially with some contribution from the -process (slow neutron-capture process) in asymptotic giant branch stars.
The most chemically similar dwarf galaxy stream is Elqui (Ji et al., 2020a), which also has low [/Fe] at . Based on its low mean metallicity similar to Leiptr, Li et al. (2022) argued Elqui was an ultra-faint dwarf galaxy using the mass-metallicity relation (as well as the Turranburra stream). But Elqui has sharply rising [Ba/Fe] with increasing metallicity and solar [Sr/Fe], which is more similar to the faintest classical dwarf spheroidal galaxies like Draco, while Leiptr has constant or declining [Sr/Fe] and [Ba/Fe] with metallicity, which is most similar to the ultra-faint dwarf galaxies. This suggests Leiptr is the stream of an even lower-mass dwarf galaxy than Elqui.
@@ 5.1 Extragalactic Origin
Leiptr’s eccentric and retrograde orbit is indicative of the progenitor’s ex-situ formation (Ibata et al., 2019, 2024). The orbit has most recently been determined by Malhan et al. (2022) to have a pericenter of 12.3 0.1 kpc and the apocenter to be 45.1 0.2 kpc, with a resulting eccentricity of 0.57. The stream is also the second-most retrograde system in that sample. Accreted systems can sometimes be linked together by looking at their energy and angular momentum, indicating that the systems have fallen into the Milky Way together, perhaps having formed with each other. Bonaca et al. (2021) suggested Leiptr could be associated with other retrograde streams, including GD-1, Phlegethon, Ylgr, and especially Gjöll and Wambelong. They suggest this group of streams originated from the retrograde Sequoia/I’itoi group. While they did not explicitly address Leiptr, Malhan et al. (2022) also associated these retrograde streams with the Sequoia/I’itoi group222The more metal-rich Arjuna system is also kinematically associated with this group (Naidu et al., 2020), but it is likely debris from an early orbit of GSE (Naidu et al., 2021).. We note that Leiptr has clearly distinct chemical abundances from Sequoia (Matsuno et al., 2022a), with similarly low [/Fe] but at lower-metallicity, , compared to Sequoia at and with lower strontium and barium abundances. Therefore, it is unlikely that Leiptr originated from a globular cluster in this dwarf galaxy as suggested as one possibility in Bonaca et al. (2021).
@@ 5.3 Low-Mass, Possibly Ultra-Faint
The strongest progenitor mass constraint is from the neutron-capture abundances derived in the Leiptr stars. Ultra-faint dwarf galaxies have , meaning that a typical ultra-faint dwarf galaxy is enriched by fewer than one rare and prolific -process event, which typically occurs once per of stars formed (Ji et al., 2016a). This explains the empirical result that the vast majority () of ultra-faint dwarf galaxies have such low strontium, barium, and europium abundances (; Ji et al. 2019; Frebel & Ji 2023) that can be used as a chemical tag to identify stars from these low-mass galaxies (e.g., Casey & Schlaufman, 2017; Roederer, 2017). Figures 11 and 12 highlight that the differences are largest at , where all dwarf galaxies with have high abundances of strontium and barium, while those with , like Boötes I and Carina II, keep low strontium and barium abundances below . Leiptr’s neutron-capture abundances are also lower than the previous lowest-mass dwarf galaxy stream documented, Elqui, whose chemical evolution is similar to the dwarf galaxy Draco (Ji et al., 2020a), suggesting Elqui had a more massive progenitor than Leiptr despite their similar mean metallicity. This suggests that Leiptr was sufficiently low-mass to still be in the stochastic neutron-capture element regime, or stellar mass .
@@ 5.4 Inhomogeneous Core-Collapse Supernova Enrichment?
We suggest Leiptr-321 may show a signature of inhomogeneous enrichment from a core-collapse supernova. Inhomogeneous Type Ia supernova enrichment has previously been suggested as a way to explain unusual magnesium and calcium scatter in Carina (Venn et al., 2012) and an “iron-rich” star in Ursa Minor (McWilliam et al., 2018). In contrast, in Leiptr we see an enhancement of both and odd-Z elements, which are synthesized by core-collapse supernovae. If this is inhomogeneous core-collapse supernova enrichment, it would allow probing the nucleosynthesis output of a single core-collapse supernova in the same manner that McWilliam et al. (2018) did for a Type Ia supernova. However, with just four member stars, we believe more observations are needed in Leiptr before such an exploration is warranted.
An alternative explanation is that the increase in and odd-Z elements is due to a late-time starburst. Such a late-time starburst has been suggested to explain a slight [/Fe] increase in the highest metallicity stars in the Magellanic Clouds (Nidever et al., 2020; Hasselquist et al., 2021). We consider this a less likely scenario than inhomogeneous enrichment, as it seems like most stars should then be in this higher- starburst component. Again, more stars in Leiptr are needed to make any conclusions.
@@ 6 Conclusion
Extragalactic Origin: Leiptr’s extremely retrograde and eccentric orbit makes it clear the progenitor formed ex-situ and was later accreted onto the Milky Way. The stars’ low [Mg/Fe] reinforce this, as they are significantly lower than the Milky Way halo stars and globular clusters (see Figure 8).
@@
Inhomogeneous Mixing?: We find an intriguing and unusual enrichment in the highest-metallicity star, Leiptr-321, which shows a significant increase in and odd-Z elements (Figures 5, 8, and 9). This abundance difference could occur due to inhomogeneously mixed material from an individual core-collapse supernova.
@@ Acknowledgments
This paper includes data gathered with the 6.5 m Magellan Telescopes, located at Las Campanas Observatory, Chile. We thank Joss Bland-Hawthorn and other members of the S5 collaboration for helpful discussions. KRA also thanks Donald Terndrup for thoughtful feedback on an earlier version of this paper submitted to the Ohio State University Libraries’ Knowledge Bank Repository. KRA acknowledges this material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2234693. SAU acknowledges support from the American Association of University Women through the American Dissertation Fellowship. SAU and APJ acknowledge support from the National Science Foundation under grants AST-2206264 and AST-2307599. TTH acknowledges support from the Swedish Research Council (VR 2021-05556). GFL, SLM, and DBZ acknowledge support from the Australian Research Council through Discovery Project grant DP220102254.

### v2

@@ 4.6 Neutron-capture elements (Sr, Ba, Eu)
Figure 11 shows that Leiptr has very low neutron-capture abundances, [Sr/Fe] [Ba/Fe] . The NLTE correction is negligible for strontium (Mashonkina & Gehren, 2001; Bergemann et al., 2012b), barium (Mashonkina et al., 1999; Mashonkina & Gehren, 2001), and europium (Mashonkina & Gehren, 2001; Guo et al., 2025). The four Leiptr member stars have significantly lower [Sr/Fe] and [Ba/Fe] than halo stars, most dwarf galaxy stars, or any of the Milky Way globular clusters. For example, globular clusters M92 and NGC 2419 have higher [Sr/Fe] between and 0, and Leiptr’s [Sr/Fe] only falls within the extreme tail end of the halo scatter in Figure 11. The systems with the most similar neutron-capture element abundances, especially for strontium, are the ultra-faint dwarf galaxies with , which typically have low [Sr/Fe] and [Ba/Fe] between and (Frebel & Norris, 2015; Kirby et al., 2017; Ji et al., 2019).
We detect [Eu/Fe] in the two cooler Leiptr members at quite low values, . These two stars have [Ba/Eu] compared to a pure heavy -process ratio of (e.g., Simmerer et al., 2004), suggesting the neutron-capture element abundances in Leiptr are dominated by production via the -process but potentially with some contribution from the -process (slow neutron-capture process) in asymptotic giant branch stars.
The most chemically similar dwarf galaxy stream is Elqui (Ji et al., 2020a), which also has low [/Fe] at .
Based on its low mean metallicity similar to Leiptr, Li et al. (2022) argued Elqui was an ultra-faint dwarf galaxy using the mass-metallicity relation (as well as the Turranburra stream).
But Elqui has sharply rising [Ba/Fe] with increasing metallicity and solar [Sr/Fe], which is more similar to the faintest classical dwarf spheroidal galaxies like Draco, while Leiptr has constant or declining [Sr/Fe] and [Ba/Fe] with metallicity, which is most similar to the ultra-faint dwarf galaxies.
This suggests Leiptr is the stream of an even lower-mass dwarf galaxy than Elqui.
@@ 5.1 Extragalactic Origin
Leiptr’s eccentric and retrograde orbit is indicative of the progenitor’s ex-situ formation (Ibata et al., 2019, 2024). The orbit has most recently been determined by Malhan et al. (2022) to have a pericenter of 12.3 0.1 kpc and the apocenter to be 45.1 0.2 kpc, with a resulting eccentricity of 0.57. The stream is also the second-most retrograde system in that sample. Accreted systems can sometimes be linked together by looking at their energy and angular momentum, indicating that the systems have fallen into the Milky Way together, perhaps having formed with each other. Bonaca et al. (2021) suggested Leiptr could be associated with other retrograde streams, including GD-1, Phlegethon, Ylgr, and especially Gjöll and Wambelong. They suggest this group of streams originated from the retrograde Sequoia/I’itoi group using the average isochrone [Fe/H] = from Ibata et al. (2019). While they did not explicitly address Leiptr, Malhan et al. (2022) also associated these retrograde streams with the Sequoia/I’itoi group222The more metal-rich Arjuna system is also kinematically associated with this group (Naidu et al., 2020), but it is likely debris from an early orbit of GSE (Naidu et al., 2021).. Given the updated metallicity from this work, we further note that Leiptr has clearly distinct chemical abundances from Sequoia (Matsuno et al., 2022a), with similarly low [/Fe] but at lower-metallicity, , compared to Sequoia at and with lower strontium and barium abundances. Therefore, it is unlikely that Leiptr originated from a globular cluster in this dwarf galaxy. Instead it is expected to be the last stream of a fully disrupted object.
@@ 5.3 Low-Mass, Possibly Ultra-Faint
The strongest progenitor mass constraint is from the neutron-capture abundances derived in the Leiptr stars. Ultra-faint dwarf galaxies have , meaning that a typical ultra-faint dwarf galaxy is enriched by fewer than one rare and prolific -process event such as a binary neutron star merger, which typically occurs once per of stars formed (Ji et al., 2016a). This explains the empirical result that the vast majority () of ultra-faint dwarf galaxies have such low strontium, barium, and europium abundances (; Ji et al. 2019; Frebel & Ji 2023) that can be used as a chemical tag to identify stars from these low-mass galaxies (e.g., Casey & Schlaufman, 2017; Roederer, 2017). Figures 11 and 12 highlight that the differences are largest at , where all dwarf galaxies with have high abundances of strontium and barium, while those with , like Boötes I and Carina II, keep low strontium and barium abundances below . Leiptr’s neutron-capture abundances are also lower than the previous lowest-mass dwarf galaxy stream documented, Elqui, whose chemical evolution is similar to the dwarf galaxy Draco (Ji et al., 2020a), suggesting Elqui had a more massive progenitor than Leiptr despite their similar mean metallicity. This suggests that Leiptr was sufficiently low-mass to still be in the stochastic neutron-capture element regime, or stellar mass . We note that the origin of the low but non-zero neutron-capture element abundances in most ultra-faint dwarf galaxies remains unknown, though it could be attributed to neutrino-driven winds, magnetorotationally driven jets, spinstars, or other unknown low-yield -process sources (Ji et al., 2019).
@@ 5.4 Inhomogeneous Core-Collapse Supernova Enrichment?
We suggest Leiptr-321 may show a signature of inhomogeneous enrichment from a core-collapse supernova. Inhomogeneous Type Ia supernova enrichment has previously been suggested as a way to explain unusual magnesium and calcium scatter in Carina (Venn et al., 2012) and an “iron-rich” star in Ursa Minor (McWilliam et al., 2018). In contrast, in Leiptr we see an enhancement of both and odd-Z elements, which are synthesized by core-collapse supernovae. If this is inhomogeneous core-collapse supernova enrichment, it would allow probing the nucleosynthesis output of material dominated by a single core-collapse supernova in the same manner that McWilliam et al. (2018) did for a Type Ia supernova. However, with just four member stars, we believe more observations are needed in Leiptr before such an exploration is warranted.
An alternative explanation is that the increase in and odd-Z elements is due to a late-time starburst. Spikes in star formation can cause rising [X/Fe] (Colavitti et al., 2009) and such a late-time starburst has been suggested to explain a slight [/Fe] increase in the highest metallicity stars in the Magellanic Clouds (Nidever et al., 2020; Hasselquist et al., 2021). While certainly a valid explanation, we consider this a less likely scenario than inhomogeneous enrichment, as it seems like most stars should then be in this higher- starburst component. However, more stars in Leiptr are needed to make any clear conclusions.
@@ 6 Conclusion
Extragalactic Origin: Leiptr’s extremely retrograde and eccentric orbit makes it clear the progenitor formed ex-situ and was later accreted onto the Milky Way. The stars’ low [Mg/Fe] reinforces this, as they are significantly lower than the Milky Way halo stars and globular clusters (see Figure 8).
@@
Inhomogeneous Mixing?: We find an intriguing and unusual enrichment in the highest-metallicity star, Leiptr-321, which shows a significant increase in and odd-Z elements (Figures 5, 8, and 9). This abundance difference could occur due to inhomogeneously mixed material from an individual core-collapse supernova. However, another possible explanation is that Leiptr experienced a late-time starburst.
@@ Acknowledgments
This paper includes data gathered with the 6.5 m Magellan Telescopes, located at Las Campanas Observatory, Chile. We thank the referees for their comments, which have helped improve this paper, and Joss Bland-Hawthorn and other members of the S5 collaboration for helpful discussions. KRA also thanks Donald Terndrup for thoughtful feedback on an earlier version of this paper submitted to the Ohio State University Libraries’ Knowledge Bank Repository. KRA acknowledges this material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2234693. SAU acknowledges support from the American Association of University Women through the American Dissertation Fellowship. SAU and APJ acknowledge support from the National Science Foundation under grants AST-2206264 and AST-2307599. TTH acknowledges support from the Swedish Research Council (VR 2021-05556). GFL, SLM, and DBZ acknowledge support from the Australian Research Council through Discovery Project grant DP220102254.